Solve by Factoring: 2x2-9x + 4 = 0
Answer #1

Answer :

Consider the given quadratic equation

2x2 - 9x + 4 = 0

2x2 - x - 8x + 4 = 0

x(2x - 1) - 4(2x - 1) = 0

(x - 4)(2x - 1) = 0

This product becomes zero means that

x - 4 = 0 , 2x - 1 = 0

x = 4 , x = 1/2

verification :

Plug in x = 4 in 2x2 - 9x + 4

2(4)2 - 9(4) + 4 = 0

so , x = 4 is a solution of the equation 2x2 - 9x + 4 = 0

Plug in x = 1/2 in 2x2 - 9x + 4

2(1/2)2 - 9(1/2) + 4 =0

so , x = 1/2 is a solution of the equation 2x2 - 9x + 4 = 0